OpenStudy (anonymous):

Simplify the expression x^-4y^3x^4y^2

OpenStudy (anonymous):

y^5

OpenStudy (anonymous):

thanks

OpenStudy (solomonzelman):

lets RVW some pro[perties of exponents, but before we do this, you know that what your equation really says is: \(\large\color{black}{ \displaystyle x^{-4} \color{red}{\times}y^3\color{red}{\times}x^4\color{red}{\times}y^2 }\) (multiplication) And also that this can be re-written as: \(\large\color{black}{ \displaystyle x^{-4} \color{red}{\times} x^4\color{red}{\times} y^3\color{red}{\times} y^2 }\)

OpenStudy (solomonzelman):

The rules of exponents say that: \(\large\color{black}{ \displaystyle a^b\times a^c=a^{b+c} }\) And that, \(\large\color{black}{ \displaystyle a^{0}=1\quad \quad \quad \color{blue}{ {\rm provided~that}~~a\ne 0}}\)

OpenStudy (solomonzelman):

So you would end up getting: \(\large\color{black}{ \displaystyle x^{-4} \color{red}{\times} x^4\color{red}{\times} y^3\color{red}{\times} y^2 }\) \(\large\color{black}{ \displaystyle x^{-4+4}\color{red}{\times} y^{3+2}=x^0\color{red}{\times}y^5=1\color{red}{\times}y^5=y^5 }\)